OpenStudy (anonymous):

integrate (5x^2+20x+6) / (x^3+2x^2+x). show steps.

OpenStudy (australopithecus):

I would just factor first than use partial fractions

OpenStudy (australopithecus):

then*

OpenStudy (australopithecus):

this requires the "harder" method of partial fractions

OpenStudy (anonymous):

yes i would do that too, and i would cheat to do it

OpenStudy (anonymous):

because computing these partial fractions is going to be a bit of a pain since the denominator is \(x(x+1)^2\) fortunately it is cooked up to give you integer answers

OpenStudy (anonymous):

@eliassaab has a snap method for doing these but i don't know it

OpenStudy (anonymous):

A = 6 5 = A + B + C 20 = 2A + B +C What am i suppose to do with that? I can't seem to solve for B or C.

OpenStudy (anonymous):

well if you know A = 6 then you have two equations with two unknowns

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i need a value for B and C

OpenStudy (anonymous):

hold on one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

assuming your equations are right, the first one becomes \[5 = 6 + B + C\] or \[-1=A+C\]

OpenStudy (anonymous):

ok something is messed up let me do it with pencil

OpenStudy (anonymous):

is there a good digital pen/pad i could buy? i wish i could show you my whole paper

OpenStudy (anonymous):

i don't know of one first off, lets get the answer

OpenStudy (anonymous):

it is \[\frac{6}{x}-\frac{1}{x+1}+\frac{9}{(x+1)^2}\]

OpenStudy (anonymous):

now we have \[\frac{5x^2+20x+6}{x(x+1)^2}=\frac{A}{x}+\frac{B}{x+1}+\frac{Cx+D}{(x+1)^2}\]

OpenStudy (anonymous):

oh wait, i didnt do it exactly like that

OpenStudy (anonymous):

here are the partial fractions, you don't have to ask wolfram to integrate http://www.wolframalpha.com/input/?i=%285x^2%2B20x%2B6%29+%2F+%28x^3%2B2x^2%2Bx%29 it does it for you

OpenStudy (anonymous):

i have x(x+1)(x+1) in denominator

OpenStudy (anonymous):

http://www.wolframalpha.com/inputwe get \[a(x+1)^2+bx(x+1)+(cx+d)x=5x^2+20x+6\]

OpenStudy (anonymous):

replacing \(x=0\) you get \(a=6\) right away

OpenStudy (anonymous):

i have A/x + B/x+1 + C/x+1

OpenStudy (anonymous):

that is a mistake

OpenStudy (anonymous):

you have a repeated factor in the denominator

OpenStudy (anonymous):

so it should be what i wrote above

OpenStudy (anonymous):

\[\frac{5x^2+20x+6}{x(x+1)^2}=\frac{A}{x}+\frac{B}{x+1}+\frac{Cx+D}{(x+1)^2}\]

OpenStudy (anonymous):

if you have a repeated factor, you need to take successive powers

OpenStudy (anonymous):

however, i too made a mistake since they are repeated it should be \[\frac{5x^2+20x+6}{x(x+1)^2}=\frac{A}{x}+\frac{B}{x+1}+\frac{C}{(x+1)^2}\]

OpenStudy (anonymous):

this gives \[a(x+1)^2+bx(x+1)+cx=5x^2+20x+6\] and as before replacing \(x\) by 0 gives \(a=6\)

OpenStudy (anonymous):

then replacing \(x\) by -1 gives \(-c=5-20+6)\) and so \(c=9\)

OpenStudy (anonymous):

i have a monstrosity in my numerator

OpenStudy (anonymous):

and since if you expand on the right you will get the coefficients of the \(x^2\) term as \(a+b\) and as \(a=6\) this is \(6+b\) you know \(6+b=5\) and so \(b=-1\)

OpenStudy (anonymous):

don't expand

OpenStudy (anonymous):

it is too much work

OpenStudy (anonymous):

ok, then A(x+1)^3

OpenStudy (anonymous):

numerator is \[a(x+1)^2+bx(x+1)+cx=5x^2+20x+6\] there is no cube term

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

start with \[\frac{5x^2+20x+6}{x(x+1)^2}=\frac{A}{x}+\frac{B}{x+1}+\frac{C}{(x+1)^2}\][

OpenStudy (anonymous):

since \(x+1\) is a repeated factor

OpenStudy (anonymous):

then if you add the stuff on the right, the numerator will be \[a(x+1)^2+bx(x+1)+cx=5x^2+20x+6\]

OpenStudy (anonymous):

k, u lost me with that step...

OpenStudy (anonymous):

or if you prefer \[A(x+1)^2+Bx(x+1)+Cx=5x^2+20x+6\] are these steps clear?

OpenStudy (anonymous):

oh ok i thought maybe so

OpenStudy (anonymous):

i thought A gets the denominators it doesn't have

OpenStudy (anonymous):

lets go back to \[\frac{5x^2+20x+6}{x(x+1)^2}=\frac{A}{x}+\frac{B}{x+1}+\frac{C}{(x+1)^2}\]

OpenStudy (anonymous):

which would be x+1 and (x+1)^2

OpenStudy (anonymous):

i see the problem

OpenStudy (anonymous):

k, ya, i agree with that step

OpenStudy (anonymous):

the denominator of your original thing is \(x(x+1)^2\), that is the least common multiple of the denominators on the right hand side as well

OpenStudy (anonymous):

so if you were going to add \[\frac{A}{x}+\frac{B}{x+1}+\frac{C}{(x+1)^2}\] then you would use your original denominator (that is the point of this)

OpenStudy (anonymous):

since the lcd is \(x(x+1)^2\) and you have the term \(\frac{A}{x}\) you only need to multiply it top and bottom by \((x+1)^2\)

OpenStudy (anonymous):

in other words when you add, that term would look like \[\frac{A(x+1)^2}{x(x+1)^2}\]

OpenStudy (anonymous):

so, in the future can i just look at the left denom and ask myself what's missing

OpenStudy (anonymous):

oh yes exaclty!

OpenStudy (anonymous):

just like you would with numbers in fact

OpenStudy (anonymous):

assuming i didnt screw up that denom

OpenStudy (anonymous):

right

OpenStudy (anonymous):

cause if it's not in LCD it wont work i assume

OpenStudy (anonymous):

so with that in mind, hopefully this line \[A(x+1)^2+Bx(x+1)+Cx=5x^2+20x+6\] makes sense

OpenStudy (anonymous):

you are thinking too hard i believe you started with the denominator, that is what you have to end up with they have to be the same

OpenStudy (anonymous):

so if we get to the point where \[A(x+1)^2+Bx(x+1)+Cx=5x^2+20x+6\] makes sense, then finding A, B and C is not that hard

OpenStudy (anonymous):

ok, ya, i'm with you now

OpenStudy (anonymous):

don't expand! it is a pain to do it that way

OpenStudy (anonymous):

lemme see if i can continue

OpenStudy (anonymous):

you can of course if you like you can expand the left hand side, then equate like coefficients, and then solve a 3 by 3 system of equations, but that is really the long way to do it substitute appropriate values for \(x\) to get what you want

OpenStudy (anonymous):

what's the better way to do it

OpenStudy (anonymous):

i usually expand the right side...

OpenStudy (anonymous):

replace \(x=0\) solve for A in one step with your eyes

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

ok, yes, A=6

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

replace \(x\) by what will give you zero in every other slot

OpenStudy (anonymous):

i.e.\(-1\) this takes only a tiny bit of arithmetic, although it is not as snappy as finding A

OpenStudy (anonymous):

is that step clear?

OpenStudy (anonymous):

no, can u give me the equation for B

OpenStudy (anonymous):

replace \(x=-1\) and you will find C right away

OpenStudy (anonymous):

on the left you will get \(-C\) and on the right you will get \(5(-1)^2+20\times -1+6=-9\) which tells you \(C=9\)

OpenStudy (anonymous):

now we can find B

OpenStudy (anonymous):

wait... how do u know to use -1.. i wouldnt see that

OpenStudy (anonymous):

because each term except \(Cx\) contains a factor of \(x+1\)

OpenStudy (anonymous):

cause i'm looking at 5x^2 + 20x + 6 = .. all that stuff in the numerator

OpenStudy (anonymous):

so if you replace \(x\) by -1 you will only be left with \(-C\) on the left

OpenStudy (anonymous):